# Why Are Linear RNNs More Parallelizable?

### 04 Mar 2026 v1

@@ C.4.4 Simulation theorem: a 44-layer DeltaNet simulates any WFA
Fix $\mathbb{K}=\mathbb{Q}$. Let $\mathcal{A}=(Q,\Sigma,\{M_{\sigma}\},\alpha,\omega)$ be an $n$-state WFA over $\mathbb{K}$ (Definition 13), and let $w=w_{1}\cdots w_{T}$ be an input word. Then there exists a 44-layer stacked DeltaNet whose scalar output at every position $t$ equals
@@ C.5 Iterated $3\times 3$ Matrix Multiplication with DeltaNet
Fix $\mathbb{K}=\mathbb{Q}$. There exists a 44-layer stacked DeltaNet such that, given the stream encoding $A^{(1)},\dots,A^{(N)}\in\mathbb{Q}^{3\times 3}$ (Definition 14), the network outputs the 99 entries of
@@
Step 1: reduce right-multiplication to a 99-dimensional linear update.
@@
Thus it suffices to maintain a 99-dimensional row state and repeatedly apply the linear map $x\mapsto xB(A)$.
@@
We will apply Lemma 12 with $n=9$, i.e. we maintain an arithmetic state of dimension $2n+1=19$ (main 99, scratch 99, temp 11), and can apply any $9\times 9$ matrix in
@@
Each input matrix $A^{(t)}$ arrives as 99 tokens. We group matrices into superblocks of length
@@
By Lemma 12 (with $n=9$), for each full superblock product $\Pi_{b}$ with $b<B$ there exists a length-$694$ sequence of $19\times 19$ matrices of the form $H(\beta,k)$. Extend this to a length-$702$ sequence by appending 88 identity steps. Thus each full superblock $b<B$ is associated with a fixed sequence
@@
Extend this to a length-$702$ sequence by appending 88 identity steps. Thus each superblock $b$ is associated with a fixed sequence
@@
Layer 4 (arithmetic in dimension 1919). Layer 4 uses one DeltaNet head of dimension 1919 and interprets the first row of $S_{t}$ as the arithmetic row vector $x\mid s\mid t]$, where $x,s\in\mathbb{Q}^{1\times 9}$ and $t\in\mathbb{Q}$.
@@
i.e. the 99 entries of $P_{N}$ in row-major order, at the final position $T=9N$. ∎
@@ Appendix G Additional Details for Experiments
Deterministic graph reachability. We construct a synthetic reachability dataset over directed graphs with $n$ vertices. Vertices are assigned to two buckets under label-dependent constraints: for positive samples, vertices 0 and $n$ lie in the same bucket, while for negative samples, they lie in different buckets. All remaining vertices are assigned independently with probability pp. Directed edges are then added deterministically by connecting consecutive vertices within each bucket, yielding two disjoint directed paths. Each instance encodes a reachability query from vertex 0 to vertex $n$ as a binary-serialized sequence.
Iterated matrix multiplication with modulus. This task requires computing prefix products of a sequence of $3\times 3$ matrices over the finite ring $\mathbb{Z}_{m}$, where $m$ is a fixed prime. Each example has the form SOS $T\mid m\mid q_{k}\mid M_{1}\mid\cdots\mid M_{T}$ EOS, where $T$ denotes the sequence length, $q_{k}$ is a fixed query index, and each matrix $M_{t}\in\{-1,0,1\}^{3\times 3}$ is sampled independently and rejected until invertible modulo $m$. Let $P_{t}=M_{1}\cdots M_{t}\bmod m$ denote the prefix product at step $t$. The target sequence is $v_{1}\mid\cdots\mid v_{T}$, where $v_{t}$ equals the $q_{k}$-th entry of $P_{t}$ modulo $m$. Models are trained with a stepwise classification loss to predict $v_{t}$ at each step. The minimum sequence length for this task is 11.
Iterated matrix multiplication without modulus. In this task, models multiply a sequence of $3\times 3$ integer matrices over $\mathbb{Z}$ and predict a single binary label. Each example has the form SOS $T\mid M_{1}\mid\cdots\mid M_{T}$ EOS, where $T$ is the number of matrices and each $M_{t}\in\{-1,0,1\}^{3\times 3}$ is sampled i.i.d. with probabilities $(0.45,0.10,0.45)$ for entries (1,0,1)(-1,0,1). Let $P_{T}=M_{1}M_{2}\cdots M_{T}$ denote the product over $\mathbb{Z}$; the target label is defined as $y=1$ iff $(P_{T})_{0,0}=0$, and $y=0$ otherwise. We optionally construct balanced splits via rejection sampling to mitigate length-dependent artifacts. Since integer products may overflow for large $T$, labeling can be performed with an optional clipping cap that truncates intermediate values during multiplication. The minimum sequence length of this task is 11.
@@ Appendix H Lower bound for $\mathsf{P}$-completeness for nonlinear input-dependent ReLU RNN
Our reduction goes from a fixed one-state Cost Register Automaton (CRA) AA over the semiring $(\mathbb{N},+,\max)$, which takes an input string $w\in\Sigma^{*}$ and accepts/rejects it. (Allender et al., 2017) constructed such an automaton such that checking whether $w$ is accepted by AA is $\mathsf{P}$-complete.
@@
The language $L(A)$ of AA is defined to be the set of $w\in\Sigma^{*}$ such that $\delta(w,\bar{0})=(q_{1},\ldots,q_{r})$ with $q_{1}>0$. That is, starting by initializing all variables to 0, we execute instructions as dictated by $w$ and accept if the final value of $x_{1}$ is positive.
@@
Given an sCRA $A=(\Sigma,\rho)$ with variables $X=\{x_{1},\ldots,x_{r}\}$, we will construct a nonlinear RNN simulating AA. More precisely, the $i$th step of the RNN simulates the $i$th step of AA. The RNN uses as states $(y_{1},\ldots,y_{r})\in\mathbb{Q}^{r}$. It starts with initial value $(0,\ldots,0)$. Given an input $w=a_{1}\cdots a_{n}\in\Sigma^{*}$, we assume that we have obtained a sequence $(A_{1},b_{1}),\ldots,(A_{n},b_{n})$ of matrices $A_{i}:\mathbb{N}^{r}\to\mathbb{N}^{r}$ and vectors $b_{i}\in\mathbb{N}^{r}$. In particular, if $\rho(a_{i})$ is of the form $x_{j}:=e$, where $e$ is either $x+y$, xyx-y then $A_{i}(\bar{p})=\delta(a,\bar{p})$ and $b_{i}=\bar{0}$. If $e$ is of the for xcx-c or $x+c$, then $A_{i}$ is the identity matrix and $b_{i}=(0,\ldots,1,\ldots,0)$, where 1 occurs at position $j$. If $e$ is $\max(x_{s},x_{t})$, then we $A_{i}:\bar{x}\mapsto\bar{y}$, where $y_{j}=x_{s}-x_{t}$ and $y_{k}=x_{k}$ ($k\neq j$).

### 05 Mar 2026 v2

@@ C.4.4 Simulation theorem: a $4$-layer DeltaNet simulates any WFA
Fix $\mathbb{K}=\mathbb{Q}$. Let $\mathcal{A}=(Q,\Sigma,\{M_{\sigma}\},\alpha,\omega)$ be an $n$-state WFA over $\mathbb{K}$ (Definition 13), and let $w=w_{1}\cdots w_{T}$ be an input word. Then there exists a $4$-layer stacked DeltaNet whose scalar output at every position $t$ equals
@@ C.5 Iterated $3\times 3$ Matrix Multiplication with DeltaNet
Fix $\mathbb{K}=\mathbb{Q}$. There exists a $4$-layer stacked DeltaNet such that, given the stream encoding $A^{(1)},\dots,A^{(N)}\in\mathbb{Q}^{3\times 3}$ (Definition 14), the network outputs the $9$ entries of
@@
Step 1: reduce right-multiplication to a $9$-dimensional linear update.
@@
Thus it suffices to maintain a $9$-dimensional row state and repeatedly apply the linear map $x\mapsto xB(A)$.
@@
We will apply Lemma 12 with $n=9$, i.e. we maintain an arithmetic state of dimension $2n+1=19$ (main $9$, scratch $9$, temp $1$), and can apply any $9\times 9$ matrix in
@@
Each input matrix $A^{(t)}$ arrives as $9$ tokens. We group matrices into superblocks of length
@@
By Lemma 12 (with $n=9$), for each full superblock product $\Pi_{b}$ with $b<B$ there exists a length-$694$ sequence of $19\times 19$ matrices of the form $H(\beta,k)$. Extend this to a length-$702$ sequence by appending $8$ identity steps. Thus each full superblock $b<B$ is associated with a fixed sequence
@@
Extend this to a length-$702$ sequence by appending $8$ identity steps. Thus each superblock $b$ is associated with a fixed sequence
@@
Layer 4 (arithmetic in dimension $19$). Layer 4 uses one DeltaNet head of dimension $19$ and interprets the first row of $S_{t}$ as the arithmetic row vector $x\mid s\mid t]$, where $x,s\in\mathbb{Q}^{1\times 9}$ and $t\in\mathbb{Q}$.
@@
i.e. the $9$ entries of $P_{N}$ in row-major order, at the final position $T=9N$. ∎
@@ Appendix G Additional Details for Experiments
Deterministic graph reachability. We construct a synthetic reachability dataset over directed graphs with $n$ vertices. Vertices are assigned to two buckets under label-dependent constraints: for positive samples, vertices 0 and $n$ lie in the same bucket, while for negative samples, they lie in different buckets. All remaining vertices are assigned independently with probability $p$. Directed edges are then added deterministically by connecting consecutive vertices within each bucket, yielding two disjoint directed paths. Each instance encodes a reachability query from vertex 0 to vertex $n$ as a binary-serialized sequence.
Iterated matrix multiplication with modulus. This task requires computing prefix products of a sequence of $3\times 3$ matrices over the finite ring $\mathbb{Z}_{m}$, where $m$ is a fixed prime. Each example has the form SOS $T\mid m\mid q_{k}\mid M_{1}\mid\cdots\mid M_{T}$ EOS, where $T$ denotes the sequence length, $q_{k}$ is a fixed query index, and each matrix $M_{t}\in\{-1,0,1\}^{3\times 3}$ is sampled independently and rejected until invertible modulo $m$. Let $P_{t}=M_{1}\cdots M_{t}\bmod m$ denote the prefix product at step $t$. The target sequence is $v_{1}\mid\cdots\mid v_{T}$, where $v_{t}$ equals the $q_{k}$-th entry of $P_{t}$ modulo $m$. Models are trained with a stepwise classification loss to predict $v_{t}$ at each step. The minimum sequence length for this task is $1$.
Iterated matrix multiplication without modulus. In this task, models multiply a sequence of $3\times 3$ integer matrices over $\mathbb{Z}$ and predict a single binary label. Each example has the form SOS $T\mid M_{1}\mid\cdots\mid M_{T}$ EOS, where $T$ is the number of matrices and each $M_{t}\in\{-1,0,1\}^{3\times 3}$ is sampled i.i.d. with probabilities $(0.45,0.10,0.45)$ for entries (1,0,1)(-1,0,1). Let $P_{T}=M_{1}M_{2}\cdots M_{T}$ denote the product over $\mathbb{Z}$; the target label is defined as $y=1$ iff $(P_{T})_{0,0}=0$, and $y=0$ otherwise. We optionally construct balanced splits via rejection sampling to mitigate length-dependent artifacts. Since integer products may overflow for large $T$, labeling can be performed with an optional clipping cap that truncates intermediate values during multiplication. The minimum sequence length of this task is $1$.
@@ Appendix H Lower bound for $\mathsf{P}$-completeness for nonlinear input-dependent ReLU RNN
Our reduction goes from a fixed one-state Cost Register Automaton (CRA) $A$ over the semiring $(\mathbb{N},+,\max)$, which takes an input string $w\in\Sigma^{*}$ and accepts/rejects it. (Allender et al., 2017) constructed such an automaton such that checking whether $w$ is accepted by $A$ is $\mathsf{P}$-complete.
@@
The language $L(A)$ of $A$ is defined to be the set of $w\in\Sigma^{*}$ such that $\delta(w,\bar{0})=(q_{1},\ldots,q_{r})$ with $q_{1}>0$. That is, starting by initializing all variables to 0, we execute instructions as dictated by $w$ and accept if the final value of $x_{1}$ is positive.
@@
Given an sCRA $A=(\Sigma,\rho)$ with variables $X=\{x_{1},\ldots,x_{r}\}$, we will construct a nonlinear RNN simulating $A$. More precisely, the $i$th step of the RNN simulates the $i$th step of $A$. The RNN uses as states $(y_{1},\ldots,y_{r})\in\mathbb{Q}^{r}$. It starts with initial value $(0,\ldots,0)$. Given an input $w=a_{1}\cdots a_{n}\in\Sigma^{*}$, we assume that we have obtained a sequence $(A_{1},b_{1}),\ldots,(A_{n},b_{n})$ of matrices $A_{i}:\mathbb{N}^{r}\to\mathbb{N}^{r}$ and vectors $b_{i}\in\mathbb{N}^{r}$. In particular, if $\rho(a_{i})$ is of the form $x_{j}:=e$, where $e$ is either $x+y$, xyx-y then $A_{i}(\bar{p})=\delta(a,\bar{p})$ and $b_{i}=\bar{0}$. If $e$ is of the for xcx-c or $x+c$, then $A_{i}$ is the identity matrix and $b_{i}=(0,\ldots,1,\ldots,0)$, where 1 occurs at position $j$. If $e$ is $\max(x_{s},x_{t})$, then we $A_{i}:\bar{x}\mapsto\bar{y}$, where $y_{j}=x_{s}-x_{t}$ and $y_{k}=x_{k}$ ($k\neq j$).